OpenStudy (anonymous):

Answers: 1 2! / 4! 70 700

OpenStudy (anonymous):

http://prntscr.com/5p43gq

hartnn (hartnn):

use the property that \(\Large n! = n(n-1)!\)

hartnn (hartnn):

so, \(\Large 8! = 8\times 7!\)

hartnn (hartnn):

\(\Large 7! = 7\times 6! \\ \implies \Large 8! = 8\times 7\times 6!\)

hartnn (hartnn):

continuing this way, we get, \(\Large 8! = 8\times 7\times 6\times 5\times 4!\)

hartnn (hartnn):

this 4! in the numerator gets cancelled with one of the 4!s in the denominator

hartnn (hartnn):

\(=\large \dfrac{8\times 7\times 6\times 5}{4!}\)

hartnn (hartnn):

now you must already be knowing \(\large 4! = 4\times 3 \times 2\times 1\) plug that in and simplify :)

OpenStudy (anonymous):

is the answer 70?

hartnn (hartnn):

did you calculate it ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\(\Huge \checkmark \)

OpenStudy (anonymous):

can you help with this one too? http://prntscr.com/5p46zd

hartnn (hartnn):

same method, did you try?

OpenStudy (anonymous):

do i just use the same method?

hartnn (hartnn):

\(\Large 17! = 17\times 16\times 15\times 14\times 13!\)

hartnn (hartnn):

\(\large \dfrac{17\times 16\times 15\times 14}{4\times 3\times 2\times 1}\)

OpenStudy (anonymous):

the answer would be 2,380

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

ty!

hartnn (hartnn):

welcome ^_^